A 11 kg hanging sculpture is suspended by a 75-cm-long, 6.0 g steel wire. When the wind blows hard, the wire hums at its fundamental frequency. What is the frequency of the hum?

What is the frequency of the hum?

Express your answer with the appropriate units.
Answer #1

tenstion in string

T = mg = 11* 9.8 = 107.8 N

now Expression for fundamental frequency

(2 L f) ^2 = T / ( m/L)

(2* 0.75* f) ^2 = 107.8 / ( 0.006/ 0.75)

f = 77.39 Hz

========

Comment in case any doubt, will reply for sure...Goodluck